OpenStudy (wade123):

Use mathematical induction to prove the statement is true for all positive integers n. 6 + 12 + 18 + ... + 6n = 3n(n + 1)

OpenStudy (wade123):

is this correct? Step 1: 6 = 3(1)(1 + 1). true n= Step 2: 16 + 12 + ... + 6k + 6(k+1) = [6 + 12 + ... + 6k] + 6(k + 1) = 3k(k + 1) + 6(k + 1) = 3(k + 1) [ k + 2 ] = 3(k + 1)[(k + 1) + 1] True for k+1 statement true for all positive integers

OpenStudy (wade123):

@ganeshie8

OpenStudy (ikram002p):

step 2 let 6 + 12 + 18 + ... + 6k = 3k(k + 1) step 3 check for k+1

OpenStudy (ikram002p):

so wanna prove 6 + 12 + 18 + ... + 6(k+1) = 3(k+1)(k + 2)

OpenStudy (wade123):

mine is wrong?

OpenStudy (ikram002p):

ur is not wrong , but when u have induction u have t set the steps in order steps for induction 1_ check for n=1 2_ assume its correct for n 3_ show its correct for n+1

OpenStudy (ikram002p):

so step3 would be like show for k+1 6 + 12 + 18 + ... + 6(k+1) = 3(k+1)(k + 2) 6 + 12 + 18 + ... + 6(k+1)= 6 + 12 + 18 + ... + 6 k +6=[6 + 12 + 18 + ... + 6 k] +6 induction step :- 6 + 12 + 18 + ... + 6k = 3k(k + 1) [6 + 12 + 18 + ... + 6 k] +6 =3k(k + 1) +6 = 3(k(k+1) +2 ) =3(k+1)(k + 2)